OpenStudy (anonymous):

Mia is three years older than twice her sister Brooke’s age. The sum of their ages is less than 30. What is the greatest age Brooke could be?

OpenStudy (compassionate):

\(\Huge\bf Welcome~to~ \color{#00B4ff}{Open}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge\bf\color{#7cc517}{Study~!!}~\) Mia = 3 + 2B M + B = 30 Solve the equation by plugging in Mia's birth. 3+ 2B + B = 30 3 + 3B = 30 3B = 27 B = 9. Now plug in nine to either equation. M + B = 30 M + 9 = 30 M = 21. Now, if you plug both in to check your equation, you'll see that 21 + 9 = 30. :) Thank you for visiting OpenStudy! Come back again, I'll be glad to help you!! :)